# HDC: Hierarchical Semantic Decoding with Counting Assistance for Generalized Referring Expression Segmentation

### v1

@@ B.2 Additional Ablation Studies
As illustrated in the main paper, we hierarchically aggregate each visual-linguistic correspondence in different granularity for multi-level semantic decoding. To further prove the rationality of our design, we visualize the aggregated activation map (originally from the semantic map) at each level in Fig. 7. As observed, with progressive mutual modality complementary across granularities, the desired regions of the referent target are fully activated with the increase of the levels.
C.2 Additional Segmentation Comparisons.
@@
As illustrated in the main paper and Fig. 6, HDC can better handle the extreme challenges of GRES under multiple/single/empty target scenarios compared with the previous SOTA method ReLA [20]. Here, we incorporate more cases in Fig. 9 to demonstrate our superiority in modeling the intricate referring association between visual and linguistic.
C.3 Segmentation Results of HDC.
@@
Fig. 8 demonstrates the effectiveness of HDC in complex generalized settings, e.g., complex geometrical relationships between instances, deceptive objects, compound sentence structures, and intricate associations between referring expressions and images.
Appendix D Limitations.
@@
Benefiting from the hierarchical semantic decoding and the explicit counting ability, our method HDC sufficiently addresses the limitations of existing GRES and achieves superiority in meeting the challenges of GRES. However, there are some potential limitations. Since the referring sentences in GRES contain multiple-target expressions, the truncation of the input text into $20$ may lose detailed descriptions of some of the targets. Although we compensate it with sentence-level textual features, it still remains unsatisfactory in terms of the incompleteness of fine-grained target description. We think that how to fully utilize the textual expression in GRES is an interesting future direction.
Appendix E Error Bar.
We have retrained our model several times on gRefCOCO [20]. The results demonstrate that the randomness of the model has little effect on the performance, i.e., the standard variation is about $0.36$.
Appendix F Broader Impacts.
Malicious use of the GRES model may lead to potential negative societal impacts, including but not limited to unauthorized surveillance or privacy-infringing tracking. However, we firmly believe that the task itself is neutral with positive implications, such as image editing and human-robot interaction.
@@
Figure 8: Segmentation results of our method in generalized settings. (a) denotes the input image, and (b) showcases the segmentation results of HDC under multiple/single/non-target situations with different referring expressions.
@@
Figure 9: Segmentation results comparison in the generalized setting. (a) denotes the input image, (b) and (c) are segmentation results of ReLA and our HDC, respectively. Empty, Single, and Multiple target referent situations with different referring expressions are in top-to-bottom accordance.

### v2

@@ 8.1 Multi-granularity Mask Aggregation Visualization
As illustrated in the main paper, we hierarchically aggregate each visual-linguistic correspondence in different granularity for hierarchical semantic decoding. To further prove the rationality of our design, we visualize the aggregated activation map (originally from the semantic map) at each level in Fig. 6. As observed, with progressive mutual modality complementary across granularities, the desired regions of the referent target are fully activated with the increase of the levels.
8.2 Additional Segmentation Comparisons
@@
As illustrated in the main paper, CoHD can better handle the extreme challenges of GRES under multiple/single/empty target scenarios compared with the previous SOTA method ReLA [22]. Here, we incorporate more cases in Fig. 8 to demonstrate our superiority in modeling the intricate referring association between visual and linguistic.
8.3 Segmentation Results of CoHD
@@
Fig. 7 demonstrates the effectiveness of CoHD in complex generalized settings, e.g., complex geometrical relationships between instances, deceptive objects, compound sentence structures, and intricate associations between referring expressions and images.
9 Limitations
@@
Benefiting from the more compatible design, i.e., hierarchical semantic decoding paradigm and the explicit counting ability, our method CoHD sufficiently addresses the limitations of existing GRES and achieves superiority in meeting the challenges of GRES. However, there are some potential limitations. Since the referring sentences in GRES contain multiple-target expressions, the truncation of the input text into $20$ may lose detailed descriptions of some of the targets. Although we compensate it with sentence-level textual features, it still remains unsatisfactory in terms of the incompleteness of fine-grained target description. We believe that how to fully utilize the textual expression in GRES is an interesting future direction.
Figure 6: Visualization in Multi-granularity Mask Aggregation. (a) and (c) indicate the input image and corresponding segmentation result of our CoHD, respectively. (b) illustrates the aggregated activation map at each granularity.
@@
Figure 7: Segmentation results of CoHD in generalized settings. (a) denotes the input image, and (b) showcases the segmentation results of CoHD under multiple/single/non-target situations with different referring expressions.
@@
Figure 8: Segmentation results comparison in the generalized setting. (a) denotes the input image, (b) and (c) are segmentation results of ReLA and CoHD, respectively. Empty, Single, and Multiple target referent situations with different referring expressions are in top-to-bottom accordance.